Solve the following ordinary differential equation.

dy/dx = (x^3) sec y


1
Expert's answer
2021-09-10T08:31:25-0400

"\frac{dy}{dx}=x^3. sec\ y
\\\frac{dy}{sec \ y}=x^3dx
\\ cos \ ydy=x^3dx"

Integrating both sides, we get

"\int cos \ ydy=\int x^3dx+c
\\sin\ y=\frac{x^4}{4}+c"